# Mutual position of two smooth quadrics over finite fields

### v1

@@ 2.3. An algebraic criterion for internal/external points
Let $n$ be odd. Let $X=\{F=0\}$ be a smooth quadric over $\mathbb{F}_{q}$ given by the equation $F(x_{1},\ldots,x_{n})=\sum_{1\leq i\leq j\leq n}a_{ij}x_{i}x_{j}=0$. Let $\Delta=\det(a_{ij})$, where $a_{ij}=a_{ji}$ for $i<j$. Given a point $P=[p_{1}:\ldots:p_{n}]\in\mathbb{P}^{n-1}(\mathbb{F}_{q})$, we have:
@@
In particular, Lemma 2.4 recovers the classical algebraic criterion for the case $n=3$; this special case for irreducible conics appears in both [10, Theorem 8.3.3] and [18, Result 1].
@@
Carlitz [6] and Jung [11] defined exterior and interior points of a quadric $X=\{F=0\}$ over $\mathbb{P}^{n-1}$ in a different way. If $P=[p_{1}:\cdots:p_{n}]\in\mathbb{P}^{n-1}(\mathbb{F}_{q})$, then $P$ is exterior to $X$ if $F(p_{1},\ldots,p_{n})$ is a non-zero square in $\mathbb{F}_{q}$, and interior to $X$ if $F(p_{1},\ldots,p_{n})$ is a non-square in $\mathbb{F}_{q}$. However, as pointed out by Primrose [16] in the MathSciNet review of [11], this terminology lacks geometric meaning since it depends on the equation of $F$ (because an interior point for $F$ becomes an exterior point for $cF$ if $c$ is a non-square). On the other hand, Bruno [5, Section 9] used the statement of Lemma 2.4 as the definition of exterior and interior points but did not offer a geometric interpretation.
3. Technical lemma from character sum estimates
We will need the following general result on estimating multiplicative character sums over possibly singular spaces due to Rojas-León [17, Theorem 1.1(a)]. We state a simplified version of Rojas-León’s theorem.
@@ 3. Technical lemma from character sum estimates
Let $q\geq 7$ be an odd prime power and $n\geq 3$. Assume that $C:f=0$ and $D:g=0$ are two distinct smooth quadrics in $\mathbb{P}^{n-1}$, defined over $\mathbb{F}_{q}$. Then
@@
We justify why such $Z$ exists. The number of tangent hyperplanes defined over $\mathbb{F}_{q}$ to either $C$ or $D$ is at most $2(q^{n-2}+q^{n-3}+\cdots+1)$. Moreover, $C\cap D$ has at most $4$ distinct irreducible components, each with dimension $n-3$. The number of hyperplanes in $\mathbb{P}^{n-1}$ containing a linear component (of dimension $n-3$) is $q+1$; if the component is not linear, there is at most one hyperplane in $\mathbb{P}^{n-1}$ containing it. Thus, the number of “bad” hyperplanes is at most $2(q^{n-2}+\cdots+1)+4(q+1)$. However, the total number of hyperplanes is $q^{n-1}+q^{n-2}+\cdots+1$, which is strictly bigger than $2(q^{n-2}+\cdots+1)+4(q+1)$ when $q\geq 7$ and $n\geq 3$.
Observe that $H(\mathbf{x})=f(\mathbf{x})g(\mathbf{x})$ defines a quartic hypersurface. For one of the hypotheses, $X\cap H\cap Z=H\cap Z$ has pure codimension $2$ in $X=\mathbb{P}^{n-1}$ since a hyperplane section of a quartic is a quartic of one less dimension (hence pure codimension $2$ in $X$). The singular locus of $X\cap H\cap Z$ is equal to $C\cap D\cap Z$, which has dimension $\delta=n-4$ because $Z\cong\mathbb{P}^{n-2}$ satisfies the conditions in the previous paragraph. Let $\chi$ be the quadratic character of $\mathbb{F}_{q}$. We have
@@
4. Proof of the main result
The main result Theorem 1.3 holds for $q<7$ because we can always increase the implicit constant in the error term. Hence, we assume $q\geq 7$ for the rest of the paper. Let $C=\{f=0\}$ and $D=\{g=0\}$ denote two smooth quadrics in $\mathbb{P}^{n-1}$. Set $n=2k+1$ for $k\in\mathbb{N}$. Let $\chi$ be the quadratic character of $\mathbb{F}_{q}$, that is,
$\chi(a)=\begin{cases}1&\text{if }a\text{ is a nonzero square,}\\ -1&\text{if }a\text{ is a non-square,}\\ 0&\text{if }a=0,\end{cases}$
for each $a\in\mathbb{F}_{q}$.
Consider the set $S_{f,g}$ of all points $P$ in $\mathbb{P}^{n-1}(\mathbb{F}_{q})$ such that $P$ is external to $C=\{f=0\}$, and $P$ is internal to $D=\{g=0\}$. Let $A$ and $B$ be the discriminants of the quadratic forms $f$ and $g$, respectively. By Lemma 2.4, a point $P$ belongs to $S_{f,g}$ if and only if
@@ 4. Proof of the main result
#$\#S_{f,g}=\sum_{\mathbf{x}\in\mathbb{P}^{n-1}(\mathbb{F}_{q})}\chi(f^{2}(\mathbf{x})g^{2}(\mathbf{x}))\cdot\left(\frac{1+\chi((-1)^{k}A\cdot f(\mathbf{x}))}{2}\right)\cdot\left(\frac{1-\chi((-1)^{k}B\cdot g(\mathbf{x}))}{2}\right)$
@@
Since $C=\{f=0\}$ is smooth, we can apply Katz’s Theorem [12, Theorem 2.2] with $d=2$:
@@
Note that Katz’s Theorem is stated for character sums over affine spaces, but we can apply it to the projective space in our setting. Indeed, $\chi(f(\mathbf{x}))=\chi(f(\lambda\mathbf{x}))$ for any non-zero scalar $\lambda$ as $\deg(f)=2$ and $\chi$ is a quadratic character. We divide by $q-1$ to account for this scaling by $\lambda$.
@@
Acknowledgements
We are grateful to Semin Yoo for helpful discussions on quadratic forms.

### v2

@@ 2.3. An algebraic criterion for internal/external points
The following lemma is folklore; see [5, Proof of Proposition 6.1]. In particular, Lemma 2.4 recovers the classical algebraic criterion for the case $n=3$; this special case for irreducible conics appears in both [11, Theorem 8.3.3] and [20, Result 1]. We include a proof for the sake of completeness.
@@
Let $n$ be odd. Let $X=\{F=0\}$ be a smooth quadric over $\mathbb{F}_{q}$ given by the equation $F(x_{1},\ldots,x_{n})=\sum_{1\leq i,j\leq n}a_{ij}x_{i}x_{j}=0$ such that $a_{ij}=a_{ji}$ for $i\neq j$. Let $\Delta=\det(a_{ij})$. Given a point $P=[p_{1}:\ldots:p_{n}]\in\mathbb{P}^{n-1}(\mathbb{F}_{q})$, we have:
@@
Carlitz [7] and Jung [13] defined exterior and interior points of a quadric $X=\{F=0\}$ over $\mathbb{P}^{n-1}$ in a different way. If $P=[p_{1}:\cdots:p_{n}]\in\mathbb{P}^{n-1}(\mathbb{F}_{q})$, then $P$ is exterior to $X$ if $F(p_{1},\ldots,p_{n})$ is a non-zero square in $\mathbb{F}_{q}$, and interior to $X$ if $F(p_{1},\ldots,p_{n})$ is a non-square in $\mathbb{F}_{q}$. However, as pointed out by Primrose [18] in the MathSciNet review of [13], this terminology lacks geometric meaning since it depends on the equation of $F$ (because an interior point for $F$ becomes an exterior point for $cF$ if $c$ is a non-square). On the other hand, Bruno [6, Section 9] used the statement of Lemma 2.4 as the definition of exterior and interior points but did not offer a geometric interpretation.
3. Proof of the main result
3.1. Technical lemmas from character sum estimates
We begin with two technical lemmas involving character sum estimates. The first one is a simplified version of a well-known result due to Katz [14, Theorem 2.2] on nonsingular multiplicative character sum estimates.
Lemma 3.1 (Katz).
Let $q$ be a prime power and $\chi$ be a nontrivial multiplicative character of $\mathbb{F}_{q}$. Let $f\in\mathbb{F}_{q}[x_{1},x_{2},\ldots,x_{n}]$ be a homogeneous polynomial with degree $d\geq 1$, such that the equation $f=0$ defines a smooth hypersurface in $\mathbb{P}^{n-1}$. If $\chi^{d}$ is trivial, then
$\left|\sum_{\mathbf{x}\in\mathbb{F}_{q}^{n}}\chi(f(\mathbf{x}))\right|\leq(d-1)^{n}q^{n/2}.$
We also need the following general result on estimating multiplicative character sums over possibly singular spaces due to Rojas-León [19, Theorem 1.1(a)]. We state a simplified version of Rojas-León’s theorem.
@@ 3.1. Technical lemmas from character sum estimates
In the rest of the paper, let $q$ be an odd prime power and let $\chi$ be the quadratic character of $\mathbb{F}_{q}$, that is,
$\chi(a)=\begin{cases}1&\text{if }a\text{ is a nonzero square in $\mathbb{F}_{q}$,}\\ -1&\text{if }a\text{ is a non-square in $\mathbb{F}_{q}$,}\\ 0&\text{if }a=0,\end{cases}$
for each $a\in\mathbb{F}_{q}$.
@@
Let $q\geq 7$ be an odd prime power and $n\geq 3$ odd. Assume that $C:f=0$ and $D:g=0$ are two distinct smooth quadrics in $\mathbb{P}^{n-1}$, defined over $\mathbb{F}_{q}$. Then
@@
We justify why such $Z$ exists. The number of tangent hyperplanes defined over $\mathbb{F}_{q}$ to either $C$ or $D$ is at most $2(q^{n-2}+q^{n-3}+\cdots+1)$. Moreover, $C\cap D$ has at most $4$ distinct irreducible components, each with dimension $n-3$. The number of hyperplanes in $\mathbb{P}^{n-1}$ containing a linear component (of dimension $n-3$) is $q+1$; if the component is not linear, there is at most one hyperplane in $\mathbb{P}^{n-1}$ containing it. Thus, the number of “bad” hyperplanes is at most $2(q^{n-2}+\cdots+1)+4(q+1)$. However, the total number of hyperplanes is $q^{n-1}+q^{n-2}+\cdots+1$, which is strictly bigger than $2(q^{n-2}+\cdots+1)+4(q+1)$ when $q\geq 7$ and $n\geq 3$. In fact, when $n\geq 5$, the intersection $C\cap D$ does not contain a hyperplane component because $C$ is parabolic (the maximum projective dimension of a linear space within $C$ is $\frac{n-3}{2}<n-3$); in this case, the number of “bad” hyperplanes can be reduced from $2(q^{n-2}+\cdots+1)+4(q+1)$ to $22(q^{n-2}+\cdots+1)+$ 2. As a result, one can check that the hyperplane $Z$ exists when $q\geq 3$ and $n\geq 5$.
Observe that $H(\mathbf{x})=f(\mathbf{x})g(\mathbf{x})$ defines a quartic hypersurface. For one of the hypotheses, $X\cap H\cap Z=H\cap Z$ has pure codimension $2$ in $X=\mathbb{P}^{n-1}$ since a hyperplane section of a quartic is a quartic of one less dimension (hence pure codimension $2$ in $X$). Since the hyperplane $Z\cong\mathbb{P}^{n-2}$ does not contain any components of $C\cap D$, the dimension of $C\cap D\cap Z$ is $\delta=n-4$. Indeed, after a projective linear change of variables, the hyperplane $Z$ is defined by the equation $\{x_{n-1}=0\}$; then, $C_{0}=C\cap Z$ and $D_{0}=D\cap Z$ are two quadrics of one less dimension (with the same equations as $C$ and $D$ after substituting $x_{n-1}=0$). Since $Z$ is tangent to neither $C$ nor $D$, we have that $C_{0}$ and $D_{0}$ are smooth quadrics. Moreover, as $Z$ does not contain any irreducible component of $C\cap D$, it follows that $C_{0}\neq D_{0}$, and so $C_{0}\cap D_{0}=C\cap D\cap Z$ has the expected dimension of $n-4$.
Claim 3.4.
The singular locus of $X\cap H\cap Z$ is equal to $C\cap D\cap Z$; in particular, the dimension of the singular locus of $X\cap H\cap Z$ is $\delta=n-4$.
Proof of the claim.
By abuse of notation, we write $Z(x_{0},\ldots,x_{n-1})=0$ for the defining equation of the hyperplane $Z$. The singular locus of $X\cap H\cap Z=H\cap Z$ is the set of points $P$ in $H\cap Z$ such that the following $2\times n$ Jacobian matrix has rank less than $2$:
$\operatorname{Jac}_{P}(f(\mathbf{x})g(\mathbf{x}),Z(\mathbf{x}))=\begin{pmatrix}f(P)\frac{\partial g}{\partial x_{0}}(P)+g(P)\frac{\partial f}{\partial x_{0}}(P)&\cdots&f(P)\frac{\partial g}{\partial x_{n-1}}(P)+g(P)\frac{\partial f}{\partial x_{n-1}}(P)\\ \frac{\partial Z}{\partial x_{0}}(P)&\cdots&\frac{\partial Z}{\partial x_{n-1}}(P)\end{pmatrix}.$
Since $P\in H\cap Z$, it follows that $f(P)=0$ or $g(P)=0$. We consider the following three cases.
  1. (1)
    If $f(P)=0$ and $g(P)=0$, then the first row vanishes, and the matrix has rank $1$.
  2. (2)
    If $f(P)=0$ and $g(P)\neq 0$, then the first row is a nonzero scalar multiple of the gradient vector $\nabla f(P)=\left(\frac{\partial f}{\partial x_{0}}(P),\ldots,\frac{\partial f}{\partial x_{n-1}}(P)\right)$. In this case, the first row is not a multiple of the second row due to the condition (iii), as the hyperplane $Z$ is not tangent to the quadric $C=\{f=0\}$. Hence, the matrix has full rank (i.e., rank $2$) in this case.
  3. (3)
    If $f(P)\neq 0$ and $g(P)=0$, the same conclusion as in (2) holds by symmetry.
Therefore, the singular locus of $H\cap Z$ is precisely $C\cap D\cap Z$. ∎
By Lemma 3.2, we have
@@
3.2. Proof of Theorem 1.3
[…]
Recall that the set $S_{f,g}$ consists of all points $P$ in $\mathbb{P}^{n-1}(\mathbb{F}_{q})$ such that $P$ is external to $C=\{f=0\}$, and $P$ is internal to $D=\{g=0\}$. Let $A$ and $B$ be the discriminants of the quadratic forms $f$ and $g$, respectively. By Lemma 2.4, a point $P$ belongs to $S_{f,g}$ if and only if
@@
#$\#S_{f,g}=\sum_{\mathbf{x}\in\mathbb{P}^{n-1}(\mathbb{F}_{q})}\chi(f^{2}(\mathbf{x})g^{2}(\mathbf{x}))\cdot\left(\frac{1+\chi((-1)^{k}A\cdot f(\mathbf{x}))}{2}\right)\cdot\left(\frac{1-\chi((-1)^{k}B\cdot g(\mathbf{x}))}{2}\right).$
@@
Since $C=\{f=0\}$ is smooth, we can apply Lemma 3.1 with $d=2$:
@@
Note that Lemma 3.1 is stated for character sums over affine spaces, but we can apply it to the projective space in our setting. Indeed, $\chi(f(\mathbf{x}))=\chi(f(\lambda\mathbf{x}))$ for any non-zero scalar $\lambda$ as $\deg(f)=2$ and $\chi$ is the quadratic character. We divide by $q-1$ to account for this scaling by $\lambda$.
@@
Acknowledgments
The authors thank Seoyoung Kim and Semin Yoo for helpful discussions. The authors also thank the anonymous referees for their valuable comments and suggestions. The research of the second author was supported in part by an NSERC fellowship.